# Donghyeon Lee1, Minbyul Jeong2∗, Jaewoo Kang1,3

### v1

@@ 6.1 Chronological Categorization
We represent ChroKnowPrompt (Chronological Knowledge Prompting), an in-depth prompting of non-parametric method to elicit chronological knowledge which has gap between each time stamp, by traversing step-by-step through the surrounding time spans.
@@
In previous section, we demonstrate whether open-source and proprietary models possess specific knowledge at various time stamps. However, this does not sufficiently assess the models’ understanding of knowledge within a chronological progression. As Zhao et al. (2024) suggest, knowledge influenced heavily by temporal factors such as general domain can still vary in more stable situation like static dataset. To address this issue, we first reclassify the models’ responses using a refined categorization scheme, allowing for a more comprehensive evaluation of temporal knowledge across all years.
Figure 4 illustrates how it works: (1) We refer it as Known for the precise temporal alignment if the model correctly identifies all relevant objects for a given knowledge category at each specific year; (2) If model fails to match the correct objects for every year, we refer it as Unknown for indicating incomplete or misaligned temporal knowledge; (3) The model accurately responds either just before or after a specific year but fails for others, signifying outdated information or forgotten legacy knowledge due to continuous updates (Cut-off); (4) The model correctly identifies some objects for a given year but not others, reflecting an incomplete understanding of the temporal knowledge (Partial Known).
Our main focus is on the Partial Known category, where models demonstrate substantial temporal knowledge but fail to answer correctly for all years, often showing confusion between correct answers. For example, Nana Akwasi Asare ($s$) was a member of sports team of ($r$) FC Utrecht ($o_{n}$) in 2011 ($t_{n}$), but the model incorrectly identifies the team as FC Groningen, despite answering correctly with FC Utrecht for 2010 ($t_{n-1}$) and in 2012 ($t_{n+1}$). At this point, we hypothesize that when the model gets one time stamp wrong, a more explicit focus on the temporal aspects surrounding that time span could help it generate more accurate answers. This is the core idea behind ChroKnowPrompt, which we will discuss details in the following section.
Figure 5: Overview of ChroKnowPrompt. The algorithm systematically traverses step by step, appending each span’s result as few shot for each steps. The range of each previous and next span is predefined, with the order of nearest time stamp from target $T_{n}$. The model suggests last candidate answer $C_{n}$, verified and refined through several steps, which ends to be checked with the object $o_{n}$ in benchmark.
@@ 6.2 Method
We introduce a chronological prompting technique for non-parametric knowledge editing, aimed at bridging knowledge gaps by utilizing multiple temporal snapshots. This method enhances the model’s reasoning by systematically integrating knowledge from different time stamps, enabling for in-depth eliciting. Figure 5 illustrates an example of the chronological prompting process.
Our process begins with a target year $t_{n}$. The algorithm systematically traverses the preceding years ($t_{n-1},t_{n-2},\ldots$) in the ‘Previous Span’ and the subsequent years ($t_{n+1},t_{n+2},\ldots$) in the ‘Next Span’. For each traversed year, the most representative object $\hat{o}_{k}$ is selected from the Correct (represented by circle) and Partial Correct (represented by triangle) categories in Table 1 using majority voting.
@@ 7 Experimental Results & Analysis
Task Configurations.
We apply our method to both Incorrect and Partial Correct categories, as the latter may still lack definitive answers. The test set consists of 10% of the total dataset from each domain. Evaluation employs fuzzy matching with a temperature of 0 for strict assessment, classifying an answer as Chrono-correct only if the last candidate answer matches the object.
@@ 7.1 Results of ChroKnowPrompt
Table 3 presents the effect of chronological prompting on knowledge elicitation across different models. Results are shown as the percentage of Known category, with parenthetical values indicating the increase by Chrono-correct. Significant improvements are observed in the biomedical domain (average increase of 11.9%), while general and legal domains show smaller gains (2.8% and 2.7%, respectively). This indicates that chronological prompting effectively enhances knowledge recall without requiring external retrieval systems. In the legal domain, models struggle to update knowledge due to the complexity of unstructured, context-rich data. Among the models, proprietary GPT4o-mini perform best overall, while open-source models like Llama-2-7b-chat-hf also show notable improvements despite outdated. These findings suggest that chronological prompting can enhance both current and legacy models by leveraging temporal context.
Table 3: Result of ChroKnowPrompt. The order of open-sources LLM is sorted by release date, starting from the latest model to the most outdated model. The numeric score is the level of Known in chronological categorization and the increase in parentheses is from the ratio of Chrono-correct which was confusing Partial correct before. Each result presents both in total span and previous span.
Models general biomedical Model Increase
total span previous span total span previous span total previous
Proprietary Large Language Models
Open-Source Large Language Models
Temporal Increase 3.2 2.4 2.2 1.3 11.4 12.3 6.2 6.0
Domain Increase 2.8 1.8 11.9 6.1
7.2 Effects of chronological span
To elucidate the mechanisms of chronological prompting, we conduct a detailed analysis across different settings, focusing on the impact of incorporating the next span in chronological contexts. Table 3 compares performance when using only the previous span versus the total span (both previous and next). Overall, the total span approach yield higher scores compared to using only the previous span. However, the degree of improvement varies by domain. In the biomedical domain, the total span results in nearly double the score increase compared to the previous span alone (11.9 vs. 6.1). Conversely, general domain shows a modest increase (2.8 vs. 1.8).
Model-specific temporal sensitivity also varies. For instance, Llama-2-7b-chat-hf demonstrates significant score increases by effectively utilizing temporal cues, particularly from next spans. On the other hand, SOLAR-10.7B-Instruct-v1.0 shows improvements when only the previous span was used, suggesting differing responses to temporal information across models. These findings indicate that models differ in their responsiveness to past and future temporal contexts. Another possible explanation is the variation in the temporal coverage of each model’s pretraining corpus, which may influence their temporal sensitivity and affect knowledge recall across different time frames.
@@ 8 Related work
8.1 Deriving and Exploring Knowledge from Language Models
@@
Since the emergence of LMs, deriving knowledge from language model is extensively studied, such as probing tasks (Hewitt & Manning, 2019), LAMA (Petroni et al., 2019) and BioLAMA (Sung et al., 2021). Then, many subsequent studies follows to exploit, (1) how LLMs define knowledge (Yu et al., ; Zhang et al., 2023; Gottesman & Geva, 2024), (2) how these models represent it (Chen et al., 2024a; b; Wang et al., 2024d), and (3) how manipulate misleading part (Wang et al., 2023; Gutiérrez et al., 2024; Wu et al., 2024a). Based on them, recent researches on investigating knowledge highlight the dynamic nature of evolving facts and suggests that contradictions within the training data may lead to knowledge conflicts (Marjanović et al., 2024; Chang et al., 2024; Wang et al., 2024a; Xu et al., 2024; Jin et al., 2024). And knowledge overshadowing (Zhang et al., 2024b) reveals phenomena where certain conditions overshadow other facts, leading to misleading information (i.e., hallucinations).
@@
In other view point, exploring temporal knowledge starts from using Wikidata, a static format of knowledge in triplet: subject, relation, and object, originated from extracting literature-based knowledge (Hahn-Powell et al., 2017). Following pioneers like TimeQA (Chen et al., 2021) and TemporalWiki (Jang et al., 2022a), many works dealing with temporal and continuous knowledge flow (Zhang & Choi, 2021; Dhingra et al., 2022; Jang et al., 2022b; Liska et al., 2022; Nylund et al., 2023; Zhu et al., 2023; Khodja et al., 2024; Zhang et al., 2024c) consist in line of it. Building upon their achievement, CarpeDiem (Kim et al., 2024) emerges to simply identify whether knowledge is outdated or not, and DyKnow (Mousavi et al., 2024) maps various models’ knowledge distribution. Also, (Zhao et al., 2024) makes dramatic work: align model into one fixed age. Though those impressive works, we try to broad the coverage of temporal knowledge. We utilize various templates to elicit the knowledge of LLMs, broaden the coverage of time stamps, and differentiate domains that should change based on a temporal perspective with those that should remain constant.
8.2 Parametric & Non-parametric Knowledge Update
@@
9 Conclusion, Limitation, and Future Work
Overall, our work highlights the critical role of temporal context in knowledge evaluation and introduces a robust framework for improving the temporal capabilities of future language models. We present ChroKnowBench, a comprehensive benchmark designed to assess the temporal knowledge of large language models (LLMs) across diverse domains, time dependencies, and temporal states. Through our novel ChroKnowledge framework, we systematically evaluate and enhance the chronological knowledge of various LLMs, revealing that the effectivenesss of knowledge elicitation depends on the specific training data formats of each model. Our findings indicate that while models often associate facts with specific time stamps, they frequently exhibit partial recall or struggle to fully capture temporal boundaries. By implementing an in-depth chronological prompting, ChroKnowPrompt, we achieve significant improvements in knowledge recall, with the most notable increase seen in the biomedical domain. This non-parametric approach proves effective for both proprietary and open-source models, demonstrating its potential as a practical method for updating and refining the temporal accuracy of LLMs without relying on external retrieval systems.
While our approach demonstrates significant improvements in certain domains, it shows limited gains in general domain. This could be a limitation of our non-parametric approach to effectively update knowledge within complex and context-rich fields. Specifically, domains like general and legal require more definitive updates that extend beyond the capabilities of prompt-based techniques alone. This underscores the need for practical knowledge updating and editing methods that are intrinsically aligned with temporal dynamics in line of parametric approach.
To address this limitation, our future work will focus on developing methods for updating the parametric knowledge of open-source LLMs. Current knowledge update techniques do not adequately handle temporal aspects, which are essential for maintaining the accuracy and relevance of information over time. We aim to create comprehensive updates synchronized with various time stamps and explore how different temporal attributes affect both structured and unstructured data. By refining the temporal alignment of knowledge updates, we hope to enhance the LLMs’ ability to recall and integrate information accurately across different time frames, thereby overcoming the current limitations observed in domains that require a deeper understanding of temporal complexity.
Appendix A Appendix
A.1 Details of Eliciting Knowledge: Few-shot Exemplars, Fuzzy Matching Rules, and Examples of Two Templates
A.1.1 Few-shot Exemplars
@@ A.1.2 Fuzzy Matching
A.1.3 Examples of Two Templates
We provide two templates of Generation and MCQA in the end of the Appendix for the better readability. For example, in Table 5, our target year is 2020 (t) to generate answer candidate of position held (r) by Donald Tusk (s).
A.1.4 Iterative Distractor Generation
@@ 8  Append $q$ to $\mathcal{H}$
A.2 Details of our source of benchmark dataset
A.2.1 Biomedical Domain

### v2

@@ 6 ChroKnowPrompt: Chronological Knowledge Prompting
Figure 5: Overview of ChroKnowPrompt. The algorithm systematically traverses step by step, appending each span’s correct answer as few shot for each steps. The range of each previous and next span is predefined, with the order of nearest time stamp from target $T_{n}$. The model suggests last candidate answer $C_{n}$, verified an d refined through several steps, which ends to be checked with the object $o_{n}$ in original ChroKnowBench.
@@ 6.1 Chronological Categorization
In previous section, we demonstrate whether open-source and proprietary models possess specific knowledge at various time stamps. However, this does not sufficiently assess the models’ understanding of knowledge within a chronological progression. As Zhao et al. (2024) suggest, knowledge influenced heavily by temporal factors as general domain can still vary in more stable situation like static dataset. To address this, we first reclassify the models’ responses using a refined categorization scheme, allowing for a more comprehensive evaluation of temporal knowledge across all years.
Figure 6 illustrates how it works: (1) Known for the precise temporal alignment if the model correctly identifies all relevant objects for a given knowledge category at each specific year; (2) If model fails to match the correct objects for every year, we refer it as Unknown for indicating incomplete or misaligned temporal knowledge; (3) The model accurately responds either just before or after a specific year but fails for others, signifying outdated information or forgotten legacy knowledge due to continuous updates (Cut-off); (4) The model correctly identifies some objects for a given year but not others, reflecting an incomplete understanding of the temporal knowledge (Partial Known).
@@
Our main focus is on the Partial Known category, where models demonstrate substantial temporal knowledge but fail to answer correctly for all years, often showing confusion between correct answers. For example, Nana Akwasi Asare ($s$) was a member of sports team of ($r$) FC Utrecht ($o_{n}$) in 2011 ($t_{n}$), but the model incorrectly identifies the team as FC Groningen, despite answering correctly with FC Utrecht for 2010 ($t_{n-1}$) and in 2012 ($t_{n+1}$). At this point, we hypothesize that when the model gets one time stamp wrong, a more explicit focus on the temporal aspects surrounding that time span could help it generate more accurate answers. This is the core idea behind ChroKnowPrompt.
@@ 6.2 Method
We introduce a chronological prompting technique for non-parametric method to elicit chronological knowledge, aimed at bridging knowledge gaps by utilizing multiple temporal snapshots. This method enhances the model’s reasoning by systematically integrating knowledge from different time stamps, enabling in-depth traverse. Our method is inspired by non-parametric editing techniques, such as Zhong et al. (2023) and Zheng et al. (2023), described in detail in Appendix A.1.
Figure 5 illustrates an example of the chronological prompting process. From a target year $t_{n}$, the algorithm systematically traverses the preceding years ($t_{n-1},t_{n-2},\ldots$) in the ‘Previous Span’ and the subsequent years ($t_{n+1},t_{n+2},\ldots$) in the ‘Next Span’. For each traversed year, the most representative object $\hat{o}_{k}$ is selected from the Correct (represented by circle) and Partial Correct (represented by triangle) categories in Table 1 using majority voting.
@@ 7.1 Results of ChroKnowPrompt
Details of task configuration is in Appendix A.8. Figure 7 presents the effect of chronological prompting on unchanged objects across different models. Results are shown as the rise of percentage in Known category with decreasing Partial Known, indicating the increase by Chrono-correct. Significant overall improvements are observed in the biomedical domain (average increase of 11.2%), while general and legal domains show smaller gains (2.6% and 3.2%, respectively). This indicates that chronological prompting effectively enhances knowledge recall without requiring external retrieval systems. In the legal domain, models struggle to recall knowledge due to the complexity of unstructured, context-rich data, especially for dynamic dataset. Among the models, proprietary GPT4o-mini and Gemini-1.5-flash perform well, while open-source models like Llama-2-7b-chat-hf also show notable improvements despite outdated.
Figure 7: Results of ChroKnowPrompt across multiple domains with unchanged objects. For each domain, the left space represents the percentage of Partial Known, and the right represents the percentage of Known. Each model includes results for both dynamic (yellow-blue bar) and static (red-green) datasets, with arrows indicating the actual increase. As shown in plots, the most effective results are observed in the biomedical domain, where the unchangeable characteristic is stronger than the general domain. While the static dataset of the legal domain shows improvement, many models struggle with unstructured format, resulting in the lowest performance among the dynamic dataset.
7.2 Analysis
Results in changed objects
Table 5 compares results for changed and unchanged objects in dynamic datasets across domains. As shown in Section 7.1, ChroKnowPrompt effectively aids models in recalling unchanged objects in both dynamic and static datasets. However, its performance on changed objects in dynamic datasets remains limited, achieving only 10–30% performance of unchanged object cases (an average of 0.43 in general domains). This highlights, despite the helpfulness and applicability of ChroKnowPrompt in addressing the chronological gaps of knowledge, recalling all historical changes of objects using prompting alone remains an exceptionally challenging problem, particularly in complex contexts such as the legal domain. These findings emphasize the need for further research into effective parametric editing techniques to assist temporal knowledge handling.
Effects of chronological span
To elucidate the mechanisms of chronological prompting, we analyze the impact of incorporating the next span in chronological contexts. As shown in Table 6, the total span (both previous and next) yields higher scores than using only the previous span with the degree of improvement varying by domain. In the biomedical domain, the total span nearly doubles the score of the previous span alone (11.9 vs. 6.1), while the general domain shows a modest increase (2.8 vs. 1.8). Model-specific temporal sensitivity also varies: Llama-2-7b-chat-hf effectively utilizes next spans, whereas Gemini-1.5-flash and SOLAR-10.7B-Instruct-v1.0 benefit more from previous spans. These suggest that differences in temporal context utilization and the coverage of pretraining corpus may influence models’ sensitivity and knowledge recall across time frames.
Effects of chat prompting
Additionally, we analyze various chat models (before instruction-tuning), including Llama-2-7b-chat-hf with findings that chronological prompting can enhance both current and legacy models by leveraging temporal context. We evaluate three open-source chat models: mpt-7b-chat (Research, 2023), Pythia-Chat-Base-7B (Biderman et al., 2023), and nemotron-3-8b-chat-4k-sft-hf (Zhang et al., 2023a). As shown in Table 6 and 7, ChroKnowPrompt is not particularly effective for chat models. Only mpt-7b-chat achieves a comparable peformance to Llama-2-7b-chat-hf (an average increase of 11.4), while Pythia-Chat-Base-7B shows almost no improvement.
@@ 8 Related work
Since the emergence of LMs, deriving knowledge from language model is extensively studied, such as probing tasks (Hewitt & Manning, 2019), LAMA (Petroni et al., 2019) and BioLAMA (Sung et al., 2021). Then, many subsequent studies follows to exploit, (1) how LLMs define knowledge (Yu et al., ; Zhang et al., 2023b; Gottesman & Geva, 2024), (2) how these models represent it (Chen et al., 2024a; b; Wang et al., 2024d), and (3) how manipulate misleading part (Wang et al., 2023; Gutiérrez et al., 2024; Wu et al., 2024a). Based on them, recent investigations of knowledge highlight the dynamic nature of evolving facts and suggests that contradictions within the training data may lead to knowledge conflicts (Marjanović et al., 2024; Chang et al., 2024; Wang et al., 2024a; Xu et al., 2024; Jin et al., 2024). And knowledge overshadowing (Zhang et al., 2024c) reveals phenomena where certain conditions overshadow other facts, leading to misleading information (i.e., hallucinations).
@@
In other view point, exploring temporal knowledge starts from using Wikidata, a static format of knowledge in triplet: subject, relation, and object, originated from extracting literature-based knowledge (Hahn-Powell et al., 2017). Following pioneers like TimeQA (Chen et al., 2021) and TemporalWiki (Jang et al., 2022a), many works dealing with temporal and continuous knowledge flow (Zhang & Choi, 2021; Dhingra et al., 2022; Jang et al., 2022b; Liska et al., 2022; Nylund et al., 2023; Zhu et al., 2023; Khodja et al., 2024; Zhang et al., 2024d) consist in line of it. Building upon their achievement, CarpeDiem (Kim et al., 2024) emerges to simply identify whether knowledge is outdated or not, and DyKnow (Mousavi et al., 2024) maps various models’ knowledge distribution. Also, (Zhao et al., 2024) makes dramatic work: align model into one fixed age. Though those impressive works, we try to broad the coverage of temporal knowledge. We utilize various templates to elicit the knowledge of LLMs, broaden the coverage of time stamps, and differentiate domains that should change based on a temporal perspective with those that should remain constant.
9 Conclusion, Limitation, and Future Work
Overall, our work highlights the critical role of temporal context in knowledge evaluation and introduces a framework for improving the temporal capabilities of future language models. We present ChroKnowBench, a benchmark for assessing temporal knowledge across diverse domains, and our ChroKnowledge framework, which evaluates LLMs’ chronological knowledge through three types of templates. Our findings indicate that while models often recall facts with time stamps, they struggle with capturing full temporal boundaries, with reliance on rigid formats like MCQA and TF. By using ChroKnowPrompt, we improve knowledge recall by reducing ambiguous Partial Known and increasing Known, particularly in biomedical domains with strong performance on unchanged objects in both proprietary and open-source models. However, challenges persist in dynamic datasets and in unstructured, context-rich formats, which amplify the difficulty of capturing temporal evolution solely with prompting. Future work will focus on the need for parametric techniques to complement prompting, enabling better alignment with changing objects and complex temporal dependencies to enhance LLMs’ temporal accuracy across various domains.
Acknowledgments
This work was supported in part by the National Research Foundation of Korea [NRF-2023R1A2C3004176, RS-2023-00262002], the Ministry of Health & Welfare, Republic of Korea [HR20C002103], and the ICT Creative Consilience program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the MSIT [IITP-2024-2020-0-01819].
Appendix A Appendix
A.1 Supplementary Studies in Knowledge Editing
A.1.1 Parametric Knowledge Update
@@ A.1.2 Non-parametric Knowledge Update
A.2 Details of Eliciting Knowledge: Few-shot Exemplars, Fuzzy Matching Rules, and Examples of Three Templates
A.2.1 Few-shot Exemplars
@@ A.2.2 Fuzzy Matching
A.2.3 Examples of Three Templates
We provide three templates of Generation, MCQA, and TF in the end of the Appendix for the better readability. For example, in Table 8 and Table 9, our target year is 2020 (t) to generate answer candidate of position held (r) by Donald Tusk (s).
A.2.4 Iterative Distractor Generation
@@ 8  Append $q$ to $\mathcal{H}$
A.3 Details of benchmark dataset
A.3.1 Statistics of Object Changes in Dynamic Dataset
The statistics of object changes among dynamic dataset in three time variant domains are in Figure 8. The figure shows how many objects have been changed among the time frame of each elements, and its distribution is proposed with the percentage of that elements across total number of dataset. The average number of object changes is 2.6 and 2.3 for general and biomedical dynamic dataset, while most of the element in legal domain has only one object changes among time frame. The biomedical domain shows the least skewness with a balanced cumulative distribution of changes, unlike the general domain, which is moderately skewed with broader change frequencies. The legal domain is highly skewed, with most changes concentrated in a single occurrence, lacking cumulative progression.
Figure 8: Statistics of object changes among dynamic dataset in three time variant domains.
A.3.2 Object-Level Focus in benchmark generation
Our ChroKnowBench emphasizes object-level changes as the core metric for assessing temporal knowledge dynamics. This approach reflects a deliberate balance between scalability, precision, and interpretability, aligning with the methodological goals of the benchmark.
Design and Scope
The benchmark evaluates how well models can align and reason about temporal knowledge by focusing on object-level transformations at specific time points. For instance: Query 1: ”In 2001, Zidane was a player at Real Madrid.” and Query 2: ”In 2010, Zidane was a coach.” By treating these as distinct evaluations, ChroKnowBench isolates object-level transitions—such as roles or affiliations—while keeping the subject and relation fixed. This design ensures a structured and scalable evaluation process that captures significant, interpretable changes in temporal knowledge.
Reason of focusing on Object-Level Changes
  • Scalability: The object-level focus simplifies the complexity of tracking temporal dynamics in subject-relation-object triples. By avoiding the combinatorial challenges associated with relational changes, this approach ensures clear and scalable evaluations.
  • Precision: Object-level changes, such as shifts in roles or affiliations, capture more nuanced updates compared to relational changes, which often default to binary states (e.g., “is a player” $\rightarrow$ “is not a player”). This granularity enhances the depth of the evaluation.
  • Flexibility: Unlike traditional Temporal Knowledge Graphs, which may impose rigid relational structures, the object-centered approach accommodates fine-grained changes in knowledge. This flexibility enables precise and interpretable assessments, particularly for dynamic, real-world transformations.
By centering evaluations on object-level changes, ChroKnowBench delivers a robust framework for measuring temporal knowledge dynamics, balancing methodological rigor with practical scalability.
A.3.3 Comparison with Temporal Knowledge Graphs
Temporal Knowledge Graphs (TKGs) are one of the well known approach for addressing temporal knowledge (Jung et al., 2020; Li et al., 2021; Zhang et al., 2024a). Our ChroKnowBench also aim to model time-sensitive knowledge, but we adopt different approaches to structuring and interpreting temporal information. In this section, we compare these two paradigms across several dimensions, with particular attention to their handling of temporal snapshots, knowledge dynamics, and suitability for domains such as biomedical data.
Temporal Snapshot vs. Knowledge-Centric Tracking
TKGs organize events based on temporal snapshots, incorporating multiple events occurring within the same timestamp into a single graph representation. This approach emphasizes the relationships between events at a specific time point, making it highly suitable for scenarios where the context of concurrent events is critical (e.g., (North America, Host a visit, Business_Africa, 2010), (Barack Obama, Consult, North America, 2010)(Zhang et al., 2024a). By connecting these events, TKGs enable reasoning about their inter-dependencies and broader temporal patterns.
In contrast, ChroKnowBench adopts a knowledge-centric perspective, focusing on the temporal evolution of individual knowledge elements (s, r, o) over time. Instead of aggregating multiple events into a single snapshot, ChroKnowBench tracks changes in object (o) values for each relation (r) across a timeline like the example in Section 3.1 This approach highlights the evolution of specific knowledge and ensures comprehensive tracking of its temporal progression.
Handling Specific Domains
A key limitation of TKGs lies in their reliance on well-defined temporal snapshots. While this approach is effective for aggregating and reasoning about concurrent events, it becomes less suitable for domains like biomedical data, where changes often unfold gradually over time and are not tied to distinct temporal snapshots. For instance, the development of a medical treatment over a decade may involve incremental advancements that cannot be neatly encapsulated within discrete, event-based snapshots.
ChroKnowBench overcomes this limitation not merely by dividing time into yearly intervals, but by prioritizing the temporal progression of individual knowledge elements. This distinction lies in its focus on tracking and organizing the dynamic and static changes of specific objects over time. While TKGs aggregate multiple concurrent events within a single snapshot, ChroKnowBench constructs yearly object pools that capture the fine-grained evolution of a specific knowledge element across its temporal trajectory. These object pools allow ChroKnowBench to explicitly track updates and fill gaps in data, ensuring a cohesive and complete representation of knowledge.
Furthermore, ChroKnowBench incorporates the concept of dynamic and static datasets, categorizing knowledge based on its temporal variability. This approach enables detailed modeling of knowledge that evolves gradually while preserving distinctions from knowledge that remains unchanged over time. By avoiding the rigid aggregation of unrelated events and instead focusing on the chronological development of individual elements, ChroKnowBench provides a more precise framework for fine-grained analysis in those specialized domains.
Comparative Summary
TKGs excel at modeling inter-event relationships within temporal snapshots, making them effective for domains where concurrent event dependencies are critical, such as geopolitical analysis. However, they struggle with gradual changes and unstructured data, limiting their applicability in multiple domains. In contrast, ChroKnowBench focuses on the detailed temporal evolution of individual knowledge units, leveraging object pools to capture gradual changes effectively, particularly in specialized domains like biomedical, and legal regulations.
Table 3 provides a comparative overview of the two paradigms.
Table 3: Comparison of TKGs and ChroKnowBench based on their temporal modeling approaches. TKGs focus on temporal snapshots aggregating multiple events, while ChroKnowBench emphasizes tracking the temporal evolution of individual knowledge elements. This table highlights their strengths, weaknesses, and domain applicability.
Aspect TKGs ChroKnowBench
Temporal Focus
Temporal snapshots aggregating
multiple events
Temporal evolution of individual
knowledge elements
Domain Applicability
Suitable for well-defined,
event-rich domains
(e.g., geopolitical, social networks)
Applicable not only temporal,
but also gradual or implicit changes
(e.g., biomedical, legal)
Handling of Gradual Changes Limited by snapshot granularity Effective through continuous tracking
Limitations
May overlook fine-grained
changes in individual knowledge
May overlook broader
event interdependencies
A.3.4 Source and Approach of Biomedical Domain